OpenStudy (anonymous):

1. x^2+8x=20 a. Add (b/2)^2 to both sides of the equation. b. Factor the left sides of the equation. c. Take the square root of both sides of the equation from step 2. d. simplify the radical and solve for x. show work.

OpenStudy (johnweldon1993):

So step a) Remember you have an equation that looks like \[ax^2 + bx = c\] So take 'b' (in your case 8) and divide it by 2 and square the result (8/2)² = 16...so we add that to both sides of the equation \[x^2 + 8x + 16 = 20 + 16\] Step b) can you factor the left hand side of the equation?

OpenStudy (johnweldon1993):

Step b) When we factor this....we see it is a perfect square ...so \[(x + 4)^2 = 36\] Step c) Take the square root of both sides...what do you get after doing this?

OpenStudy (johnweldon1993):

\[\sqrt{(x + 4)^2} = \sqrt{36}\] Simplifes to \[x + 4 = \pm6\] So now you have 2 equations \[x + 4 = 6\] and \[x + 4 = -6\] Solve both equations for 'x' and you will have your answers

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

you helped me alot

OpenStudy (johnweldon1993):

No problem! Just hope it made sense?

OpenStudy (anonymous):

it did! you should help me on other ones

OpenStudy (johnweldon1993):

lol well if I see em posted and no one else helps then I'll be happy to lol

OpenStudy (anonymous):

lol ok